OpenStudy (anonymous):

find the solution set of the equation 3(sec x-1)= 5- sec x in the interval of 0< x<2 pi. A) pi/3 and 5pi/3 B) pi/6 and 11pi/6 C) pi/6 and 5pi/6 D) pi/3 and 2pi/3

OpenStudy (anonymous):

please helo

OpenStudy (anonymous):

secx=2. cosx=1/2, x=pi/3 and 5pi/3

OpenStudy (anonymous):

@alexwee123 you add the secx in your step 3. So 4secx=8 secx=2

OpenStudy (alexwee123):

oh lol sorry about that :/

OpenStudy (alexwee123):

3(sec x-1)= 5- sec x 3secx -3=5-secx 4secx=8 secx=2 cosx=1/2 cos^-1(cosx)=cos^-1(1/2) x=pi/3+pi(n) where n=integers because range is within 0< x<2 pi. the answers are x=pi/3, 5pi/3

OpenStudy (anonymous):

thank you for your help